# Conformal prediction after efficiency-oriented model selection

### v2

@@ Appendix B Implementation for ModSel-CP-LOO
While it might appear that this has not simplified the calculation, in cases such as the residual score, these sets are more tractable to work with—specifically, for the residual score and with loss equal to interval length,
@@
Under the notation and definitions above, in the case of $\mathcal{Y}=\mathbb{R}$, assume also that $y\mapsto S^{\lambda}(x,y)$ is continuous, and $q\mapsto\mathcal{L}_{n+1}(\lambda,q)$ is continuous (e.g., as for the residual score, and for loss given by interval length). Then the partition $\mathbb{R}=\cup_{j\in\mathcal{J}}\mathcal{B}_{j}$ can be defined with the following set of breakpoints between intervals:
@@ Appendix C Proofs for theoretical results in Section 4
In this section, we will prove Theorem 4.2, which establishes that ModSel-CP and ModSel-CP-LOO achieve asymptotically optimal prediction set size under certain conditions. While Theorem 4.2 works in the special case of the residual score, $S^{\lambda}(x,y)=|y-f_{\lambda}(x)|$, here we will present a more general result showing that these optimality guarantees can hold more broadly.
@@ C.1.2 The residual score as a special case
Appendix D Additional simulations: classification
@@
Figure 5: Results for the classification experiment (see Section D for details).

### v3

@@ Appendix B Implementation for ModSel-CP-LOO
While it might appear that this has not simplified the calculation, in cases such as the residual score, these sets are more tractable to work with—specifically, for the residual score and with loss equal to interval width,
@@
Under the notation and definitions above, in the case of $\mathcal{Y}=\mathbb{R}$, assume also that $y\mapsto S^{\lambda}(x,y)$ is continuous, and $q\mapsto\mathcal{L}_{n+1}(\lambda,q)$ is continuous (e.g., as for the residual score, and for loss given by interval width). Then the partition $\mathbb{R}=\cup_{j\in\mathcal{J}}\mathcal{B}_{j}$ can be defined with the following set of breakpoints between intervals:
@@ Appendix C Proofs for theoretical results in Section 4
In this section, we will prove Theorem 4.2, which establishes that ModSel-CP and ModSel-CP-LOO achieve asymptotically optimal prediction set width under certain conditions. While Theorem 4.2 works in the special case of the residual score, $S^{\lambda}(x,y)=|y-f_{\lambda}(x)|$, here we will present a more general result showing that these optimality guarantees can hold more broadly.
@@ C.1.2 The residual score as a special case
Appendix D ModSel-CP v.s. ModSel-CP-LOO: another perspective
Since YK-baseline does not correct for selection bias, any valid method will need to be more conservative in terms of the criterion for including $y$ into the prediction set. Again, for simplicity, consider the case of the residual score, $S^{\lambda}(x,y)=|y-f_{\lambda}(x)|$, to allow for a more intuitive comparison.
Under this special score, the ModSel-CP-LOO prediction set is simplified to,
$\widehat{C}_{\texttt{ModSel-CP-LOO}}(X_{n+1})=\\ \left\{y\in\mathcal{Y}:S^{\hat{\lambda}}(X_{n+1},y)\leq\mathrm{Quantile}_{(1-\alpha)(1+1/n)}\left(S^{\hat{\lambda}_{-1}(y)}(X_{1},Y_{1}),\dots,S^{\hat{\lambda}_{-n}(y)}(X_{n},Y_{n})\right)\right\}.$
We see that ModSel-CP-LOO prediction set is constructed by adjusting the value of the RHS term (and keeping the LHS term the same). Conversely, comparing to the expression (11) in Theorem 3.2, we see that the ModSel-CP prediction set is constructed by adjusting the value of the LHS term and keeping the RHS term the same. Since two methods adjust terms differently to be more conservative so that selection bias is accounted for, we would expect the width of their resulting sets may differ under different settings.
For ModSel-CP-LOO, since it enlarges the set by enlarging the RHS term, namely, the adjusted empirical quantile (ignoring variation of model sources), then under settings where the competing models make similar predictions and yet $S^{\lambda}(X,Y)$ exhibits heavy tails, it is likely that ModSel-CP-LOO is more conservative than ModSel-CP. Evidence for this speculation can be found in the simulation results in Section 5.1, by comparing residual score results between “NormalX + sparse structure + Gaussian noise” and “tX + sparse structure + Gaussian noise”.
For ModSel-CP, on the other hand, its major source of conservativeness comes from model multiplicity, i.e., when competing models are indistinguishable via score quantile, it could possibly result in ModSel-CP returning a large set. To be more specific, by (11), if there are many competing models ($|\mathcal{M}|$ large) while each model can predict differently, the ModSel-CP set would be wide. This could be the case when both the data distribution and the model class are highly symmetrical. To demonstrate this point, we provide an additional simulation below.
  • Distribution of $(X,Y)$: $Y=X+\epsilon$, where $X\sim\mathcal{N}(0,1)$ and $\epsilon\sim\mathcal{N}(\mu,1)$. $X$ is independent of $\epsilon$.
  • Models training: For any $C>0$, the model class $\Lambda_{C}$ consists of two models:
    $\Lambda_{C}=\left\{S^{+}(x,y)=|y-x-C|,S^{-}(x,y)=|y-x+C|\right\}.$
  • As in Section 5.1, we fix $n=200$, $\alpha=0.1$. For each $C\in\{0.5,1,5\}$ and each $\mu\in\{-1,-0.5,0,0.5,1\}$, we run $5000$ independent trials of the simulation. The results are displayed below. Since YK-adjust always outputs prediction sets that are much larger than other methods, and we are comparing the performance difference between ModSel-CP and ModSel-CP-LOO under different settings, we omit displaying YK-adjust in Figure 5 to make the performance distinctions more prominent. The full results are reported in the table summary in Table 1 with standard errors in parentheses.
Figure 5: Results of additional simulation in Appendix D.
Table 1: Results of simulation in Appendix D
$C=0.5$, $n=200$
Coverage
$\mu$ ModSel-CP ModSel-CP-LOO YK-baseline YK-split YK-adjust
$-1.0$ 0.910 (0.004) 0.910 (0.004) 0.910 (0.004) 0.910 (0.004) 0.981 (0.002)
$-0.5$ 0.904 (0.004) 0.904 (0.004) 0.904 (0.004) 0.905 (0.004) 0.983 (0.002)
$0.0$ 0.893 (0.004) 0.891 (0.004) 0.879 (0.005) 0.896 (0.004) 0.976 (0.002)
$0.5$ 0.900 (0.004) 0.900 (0.004) 0.900 (0.004) 0.901 (0.004) 0.980 (0.002)
$1.0$ 0.909 (0.004) 0.909 (0.004) 0.909 (0.004) 0.908 (0.004) 0.980 (0.002)
Width of prediction interval
$\mu$ ModSel-CP ModSel-CP-LOO YK-baseline YK-split YK-adjust
$-1.0$ 3.701 (0.003) 3.701 (0.003) 3.701 (0.003) 3.721 (0.004) 5.240 (0.006)
$-0.5$ 3.312 (0.003) 3.312 (0.003) 3.312 (0.003) 3.332 (0.004) 4.729 (0.005)
$0.0$ 3.722 (0.006) 3.699 (0.004) 3.582 (0.003) 3.726 (0.005) 5.023 (0.005)
$0.5$ 3.312 (0.003) 3.312 (0.003) 3.312 (0.003) 3.337 (0.004) 4.732 (0.006)
$1.0$ 3.707 (0.003) 3.707 (0.003) 3.707 (0.003) 3.728 (0.004) 5.250 (0.006)
$C=1$, $n=200$
Coverage
$\mu$ ModSel-CP ModSel-CP-LOO YK-baseline YK-split YK-adjust
$-1.0$ 0.903 (0.004) 0.903 (0.004) 0.903 (0.004) 0.900 (0.004) 0.978 (0.002)
$-0.5$ 0.910 (0.004) 0.910 (0.004) 0.910 (0.004) 0.912 (0.004) 0.985 (0.002)
$0.0$ 0.899 (0.004) 0.899 (0.004) 0.886 (0.005) 0.898 (0.004) 0.973 (0.002)
$0.5$ 0.901 (0.004) 0.901 (0.004) 0.901 (0.004) 0.901 (0.004) 0.980 (0.002)
$1.0$ 0.895 (0.004) 0.895 (0.004) 0.895 (0.004) 0.895 (0.004) 0.980 (0.002)
Width of prediction interval
$\mu$ ModSel-CP ModSel-CP-LOO YK-baseline YK-split YK-adjust
$-1.0$ 3.313 (0.003) 3.313 (0.003) 3.313 (0.003) 3.340 (0.004) 4.736 (0.005)
$-0.5$ 3.703 (0.003) 3.703 (0.003) 3.703 (0.003) 3.724 (0.005) 5.244 (0.006)
$0.0$ 4.711 (0.011) 4.592 (0.005) 4.453 (0.003) 4.621 (0.005) 5.964 (0.005)
$0.5$ 3.703 (0.003) 3.703 (0.003) 3.703 (0.003) 3.726 (0.005) 5.245 (0.006)
$1.0$ 3.312 (0.003) 3.312 (0.003) 3.312 (0.003) 3.334 (0.004) 4.744 (0.005)
$C=5$, $n=200$
Coverage
$\mu$ ModSel-CP ModSel-CP-LOO YK-baseline YK-split YK-adjust
$-1.0$ 0.904 (0.004) 0.904 (0.004) 0.904 (0.004) 0.904 (0.004) 0.975 (0.002)
$-0.5$ 0.894 (0.004) 0.894 (0.004) 0.894 (0.004) 0.895 (0.004) 0.980 (0.002)
$0.0$ 0.892 (0.004) 0.893 (0.004) 0.880 (0.005) 0.899 (0.004) 0.970 (0.002)
$0.5$ 0.903 (0.004) 0.903 (0.004) 0.903 (0.004) 0.905 (0.004) 0.981 (0.002)
$1.0$ 0.896 (0.004) 0.896 (0.004) 0.896 (0.004) 0.897 (0.004) 0.981 (0.002)
Width of prediction interval
$\mu$ ModSel-CP ModSel-CP-LOO YK-baseline YK-split YK-adjust
$-1.0$ 10.584 (0.003) 10.584 (0.003) 10.584 (0.003) 10.609 (0.005) 12.197 (0.006)
$-0.5$ 11.582 (0.003) 11.582 (0.003) 11.582 (0.003) 11.612 (0.005) 13.196 (0.006)
$0.0$ 13.658 (0.047) 12.581 (0.005) 12.446 (0.003) 12.613 (0.005) 13.964 (0.005)
$0.5$ 11.591 (0.003) 11.591 (0.003) 11.591 (0.003) 11.620 (0.005) 13.199 (0.006)
$1.0$ 10.586 (0.003) 10.586 (0.003) 10.586 (0.003) 10.610 (0.005) 12.200 (0.006)
Appendix E Additional simulations: classification